OpenStudy (anonymous):

12^√125x^3) Simplify

OpenStudy (dumbcow):

is that a square root or cubed root ?

OpenStudy (anonymous):

square

OpenStudy (anonymous):

well the x is cubed

OpenStudy (dumbcow):

12^(5x*sqrt(5x))

OpenStudy (dumbcow):

you cant really do much with this

OpenStudy (anonymous):

its for pre calc honors summer assignment. and i need to show work >.<

OpenStudy (dumbcow):

well 125 = 5^3 5^3x^3 = 5^2x^2*5x by taking square root, 5^2x^2 becomes 5x

OpenStudy (dumbcow):

out of curiosity, whats the project about?

OpenStudy (anonymous):

its just about a whole bunch of math problems. 130 to be exact.

OpenStudy (dumbcow):

oh nevermind, that will keep you busy...

OpenStudy (anonymous):

I have most of it done, there are just a few that I have no clue how to do, about to post another one.

OpenStudy (dumbcow):

ok ill take a look